OWG Meeting 25-Mar-2015

From GlueXWiki
Jump to: navigation, search

Location and Time

Room: CC F326-327

Time: 1:30pm-2:30pm


(if problems, call phone in conference room: 757-269-6460)

  1. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 120390084.
  2. To join via a Web Browser, go to the page https://bluejeans.com/120390084.
  3. To join via phone, use one of the following numbers and the Conference ID: 120390084
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. More information on connecting to bluejeans is available.


  1. Announcements
    • .Xauthority files moved to local disks (~hdops/.Xauthority -> /tmp/.Xauthority_hdops)
    • halldchx1 printer is now accessible from gluons
    • future: implementation of local hosts file to reduce DNS queries
  2. Major systems reports
    • Trigger
    • DAQ
    • Monitoring
    • Controls
  3. hdops account policy
    • Existing policy
    • Proposed modifications to Online policy
      1. detector groups may use detector system specfic accounts for operations that might otherwise interfere with standard operations. e.g. use hdstops for running a ST specific DAQ system during long down times
      2. the hdops account will be maintained to use a standard environment for all gluon machines via the files:
        • ~hdops/.tcshrc
        • /gluex/etc/hdonline.cshrc
      3. strike the final paragraph of the 2013 policy regarding the location of COOL_HOME

Previous Meeting


Attendees: David L. (chair), Beni Z., Sergey F., Alex S., Dave A., Vardan G., Bryan M., Justin S., William G., Alex B., Simon T., Carl T.


  • Have been doing threshold scans for FCAL
    • Bases generating internal pulses. Not sure why. Adesh is looking into it.
    • Masked about 10 more noisy channels
    • Will need to run with either a high global threshold or raise the per cell threshold to take beam data
  • BCAL is running OK
    • Need to increase integration window size (only ~70% is currently being integrated)
  • Have been testing playback features and multiple triggers
    • Multiple triggers work and Alex understands them, but more timing work is needed to implement them in the standard DAQ. In particular, we discussed FCAL pulser triggers and random triggers. These will eventually be done, but are not at the top of the current priority list
  • Pair Spectrometer + TAGGER readout from standard hdops DAQ is highest priority once beam comes back


  • Running cosmics routinely without problem
  • Unable to start full DAQ in multi-event block mode (no triggers)
  • Some ROCS occasionally having problems connecting to platform
    • Seems to be correlated with ROCS that have a lot of caput calls going on (e.g. rocBCAL14 has many more than rocBCAL13 and the former seems to be running caput via the shell much more often.)


  • Work has been done to optimize the online plugins so that all but 2 run fast
  • A Secondary ET system has been implemented with an et2et program to copy events from the ER's ET system to it.
  • Separate station on the secondary ET is created for each monitoring "level" and a separate level is now used for each plugin
  • TAGM_online plugin creates more than half of the 2k histograms in the online system and is responsible for a similar share of the memory usage.
    • May need to be reduced by implementing an "online" mode for plugins that reduces number of bins and doesn't create some histograms
  • RSMonitor program written and used to profile RootSpy system
  • RunLogXXXXXX directory that is created by CODA run scripts and filled with L1 and DAQ configuration files is now tarballed automatically and only the tar ball copied to tape.
  • May do something similar for online calibration processes. David and Sean will work on this while Sean is here in in April.